Find the amount of time Joel spent watching space adventures movies.
1) He saw twice as many 1 1/2 hour space movies as he did 2-hour mysteries.
2) He spent a total of 15h watching movies.
I need to do this in a table and find the answer I am very confused.

Answer by ikleyn(52784)

.
Let "n" be the number of the 2-hour mysteries.


Then the number of the Space Adventures movies is 2n.


The total time spent equation is


    2n + 1.5*(2n) = 15  hours.


    2n + 3n = 15


    5n = 15   ====>  n = 15/5 = 3.


Answer.  Three 2-hours mysteries  and  six  1.5-hours  space movies.


         6*1.5 = 9 hours was spent for space movies.